OpenStudy (natsent06):

Derek’s favorite clothing store is having their annual fall sale. All clothing is 25 percent off. A shirt he really likes has an original price of $28. Derek has $24 in his wallet. He is wondering if he has enough money to buy the shirt now that it is on sale. He knows he must also pay sales tax. The sales tax rate is 8 percent and is calculated on the discounted price. Does Derek have enough money to buy the shirt? Justify your answer. Your explanation should contain answers to the following questions: 1. What is the amount of the discount? 2. What is the sale price? 3. How much will Derek pay in sales tax? 4. What is the total cost of the shirt? 5. Does Derek have enough money to buy the shirt?

OpenStudy (anonymous):

28 * 0.25 = ?

OpenStudy (natsent06):

7

OpenStudy (anonymous):

So it's 8 dollars off.

OpenStudy (anonymous):

7*

OpenStudy (anonymous):

lol fail.

OpenStudy (anonymous):

28 - 7 = 21

OpenStudy (anonymous):

So now we have taxes. 21 * 0.08 =

OpenStudy (natsent06):

1.68

OpenStudy (anonymous):

21 + 1.68 = 22.68

OpenStudy (anonymous):

1. What is the amount of the discount?

OpenStudy (anonymous):

We calculated it.

OpenStudy (anonymous):

You just could've searched google.... http://openstudy.com/study#/updates/531dfad1e4b0108f77dc1f8f

OpenStudy (natsent06):

so

OpenStudy (anonymous):

We calculated the discount, how much did we subtract from the cost?

OpenStudy (natsent06):

um

OpenStudy (anonymous):

28 - 7....

OpenStudy (anonymous):

7 is our discount.

OpenStudy (anonymous):

2. What is the sale price? 28 - 7

OpenStudy (natsent06):

ohhh

OpenStudy (anonymous):

3. How much will Derek pay in sales tax? 21 * 0.08 =

OpenStudy (natsent06):

oh oh 21

OpenStudy (anonymous):

4. What is the total cost of the shirt? 21 + 1.68 = 22.68

OpenStudy (natsent06):

for 3 it would be 1.68, and for 4 it would be 22.68?

OpenStudy (anonymous):

Yeah

OpenStudy (natsent06):

and he does have enough

OpenStudy (anonymous):

Unless the laws of the universe changed, he does have enough. :P

OpenStudy (natsent06):

lol

OpenStudy (natsent06):

thank you

OpenStudy (anonymous):

No problem